### v1

# New determination of the neutrino hadronic production cross sections from GeV to beyond PeV energies

Abstract
The flux of astrophysical neutrinos is now measured with unprecedented accuracy and on several decades of energy spectrum. Their origin traces back to hadronic collisions between protons and nuclei in the cosmic rays with hydrogen and helium in the target gas. To accurately interpret the data, a precise determination of the underlying cross sections is therefore mandatory. We present a new evaluation of the neutrino production cross section from $p+p$ collisions, building on our previous analysis of the production cross section for $\pi^{\pm}$, $K^{\pm}$, and minor baryonic and mesonic channels. Cross sections for scatterings involving nuclei heavier than protons are also derived. The novelty of our approach is the analytical description of the Lorentz invariant cross section $\sigma_{\rm inv}$, and the fit of the model to the available accelerator data. We work with neutrino energies from $10$ GeV to $10^{7}$ GeV, and, correspondingly, to incident proton (nuclei) energies from $10$ GeV to $10^{9}$ GeV (GeV/n). We obtain the total differential cross section, $d\sigma(p+p\rightarrow\nu+X)/dE_{\nu}$ as a function of neutrino and proton energies, with an estimated uncertainty of 5% for neutrino energies below 100 GeV, increasing to 10% above TeV energies. Predictions are given for $\nu_{e},\nu_{\mu},\bar{\nu_{e}}$ and $\bar{\nu_{\mu}}$. A comparison with state-of-the-art cross sections, all relying on Monte Carlo generators, is also presented. To facilitate the use by the community, we provide numerical tables and a script for accessing our energy-differential cross sections.
I Introduction
The main origin of astrophysical neutrinos is intimately connected to hadronic processes. Being electrically neutral, neutrinos cannot be directly accelerated by electric fields; instead, they are produced in hadronic interactions when energetic particles collide with ambient matter. Such processes occur both in extragalactic sources, such as blazars, particularly in their flaring activity periods Aartsen et al. (2018), and star forming galaxies, for example from NGC 1068 Abbasi et al. (2022). The IceCube collaboration has measured a neutrino diffuse flux of astrophysical origin in the TeV-PeV energy range and with a spectral shape compatible with a power-law of index of about $2.5$ (see, e.g., Aartsen et al. (2013, 2020); Abbasi et al. (2021, 2024)). On the other hand the detection of a 120 PeV event by KM3NeT Aiello et al. (2025) has further extended the frontier of neutrino physics. Recently, the IceCube Collaboration has found strong evidence for a Galactic plane emission, which is consistent with the diffuse neutrino flux produced from interactions of cosmic rays (CRs)—predominantly protons and helium nuclei—with the interstellar medium (ISM) via hadronic collisions Abbasi et al. (2023), even if a part of the events could arise from a population of unresolved point sources. Much like the diffuse emission of high-energy photons, neutrinos can yield valuable information on the spatial and spectral distribution of Galactic CRs throughout the Milky Way Schwefer et al. (2023). In fact, at variance with $\gamma$ rays, which can originate from both hadronic and leptonic processes, the diffuse neutrino flux provides a cleaner tracer of hadronic interactions of CRs.
Any prediction of the neutrino emission, whether Galactic or from extragalactic sources, depends on several factors, including the spectra of incident CRs, the choice of target gas maps, and the hadronic interaction cross sections. In this work, we focus specifically on improving the estimation of the latter, presenting a new model for neutrino production cross sections, building on our previous analyses reported in Orusa et al. (2022, 2023). Recently, the importance of cross sections in astroparticle physics has been extensively discussed in Maurin et al. (2025). The most direct applications of our new cross sections span processes of astrophysical neutrino production, both in diffuse or point source emissions. The applicability of our results can nonetheless be wider.
For example, the inclusive neutrino production cross sections we provide for $p+p\to\text{hadrons}\to\nu$ could be applicable to proton beam-dump and fixed-target searches for light dark matter, where neutrino-induced neutral-current interactions constitute leading backgrounds. In contrast, for central missing-transverse-momentum (MET) searches at the LHC (monojet, $\gamma+$MET, jets+MET), the dominant Standard Model backgrounds arise from electroweak processes—chiefly $pp\to Z(\nu\bar{\nu})+\text{jets}$ (irreducible) and $pp\to W(\ell\nu)+\text{jets}$ with an undetected lepton—along with top and diboson production ATLAS Collaboration (2021); Tumasyan et al. (2021), which are not the main focus of this paper. Neutrinos from pion and kaon decays produced in generic hadronic activity are predominantly forward and low transverse momentum, and therefore contribute negligibly to large MET in the central detectors. Our cross sections could therefore be directly applied to predict neutrino fluxes in forward LHC detectors and to evaluate neutrino backgrounds in proton beam-dump configurations, after appropriate generalization to $p+A$ collisions and inclusion of geometry and shielding effects. Abreu et al. (2023); Albanese et al. (2023); Batell et al. (2022); Ahdida et al. (2022).
The diffuse hadronic neutrino emission depends on the CR fluxes, the density of the ISM, and the inelastic production cross section $\sigma(p+p\rightarrow\nu+X)$ (and analogously for heavier nuclear components in both CRs and the ISM). Local CR fluxes are measured with high precision by AMS-02 Aguilar et al. (2021), DAMPE An et al. (2019), and CALET Adriani et al. (2022), and the ISM density in our nearby Galactic environment, within a few kiloparsecs, is also relatively well constrained Widmark et al. (2022). However, beyond the solar neighborhood, the situation becomes more uncertain: CR fluxes must be extrapolated from local measurements, making predictions model-dependent. Likewise, determining the distribution of gas in more distant regions of the Galaxy is considerably more challenging Pohl et al. (2008); Mertsch and Phan (2022).
A key ingredient for accurately predicting the hadronic diffuse $\nu$ emission is the inclusive neutrino production cross section $\sigma(p+p\rightarrow\nu+X)$, and other reactions involving He and heavier nuclei. The standard approach to calculate these cross sections is to employ Monte Carlo event generators Kamae et al. (2006); Kachelrieß et al. (2019); Bhatt et al. (2020); Mazziotta et al. (2016). The most widely used model is based on a customized implementation of Pythia 6 by Kamae et al. Kamae et al. (2006). For example this is the cross section theoretical framework employed by the IceCube Collaboration in Abbasi et al. (2023). More recent results include those from AAfrag Kachelrieß et al. (2019), based on the QGSJET-II-04m event generator, and from Bhatt et al. Bhatt et al. (2020), based on the DPMJET-III-19.1 event generator. Significant discrepancies between Monte Carlo simulations and experimental data have been pointed out—for example, in the production cross sections of $\bar{p}$ Kachelriess et al. (2015); Kachelrieß et al. (2020), in $e^{\pm}$ Orusa et al. (2022) (hereafter ODDK22), and in $\gamma$ rays Orusa et al. (2023). Moreover, as shown in Refs. Koldobskiy et al. (2021); Dörner et al. (2025), the predicted production cross sections of all $\nu$ flavors can differ by up to a factor of $\mathcal{O}(2)$ depending on the Monte Carlo generator used. These large variations underscore the need for improved modeling of neutrino production cross sections.
In this paper, we present a new accurate model that relies primarily on an analytic prescription fitted to the available experimental data, with the aim of determining their correct dependence on kinematic variables and of robustly quantifying the modeling uncertainties. The dominant production channel for $\nu_{e}$, $\nu_{\mu}$, $\bar{\nu_{e}}$ and $\bar{\nu_{\mu}}$ is the decay of $\pi^{\pm}$ mesons, for which we build upon the analysis performed in ODDK22. In addition, we carefully model the production cross sections of $K$ mesons and $\Lambda$ baryons, which contribute as well to the $\nu$ flux through their decay into $\pi^{\pm}$. Our strategy closely follows the approach adopted in ODDK22 for deriving cross sections relevant to the secondary production of CR electrons and positrons.
The remainder of this paper is organized as follows. In Sec. II, we outline the theoretical framework used to derive the observed $\nu$ flux from hadronic production cross sections. Section III focuses on the analytical modeling of the $\pi^{\pm}$ production cross section, the dominant channel for neutrino generation. In Sec. IV, we estimate contributions from additional production channels and from interactions involving nuclei. Our results are presented in Sec. V, followed by our conclusions in Sec. VI.
Figure 1: This diagram shows the $\nu_{\mu}$ and $\nu_{e}$ production channels from $p+p$ collisions considered in our analysis. The same scheme applies to $\bar{\nu}_{\mu}$ and $\bar{\nu}_{e}$ under charge conjugation (with the exception of the initial $p+p$ state). Only channels contributing at least 0.5% of the total yield are reported (see main text for details).
II Neutrino production cross sections and $\nu$ emissivity
The original results of this paper relate to a new determination of the $\nu$ production cross section from hadronic collisions. Throughout this section, we refer generically to the neutrino flavor as $\nu$, while the contributions from individual flavors will be discussed in more detail in the following sections. Given the relevance of these processes in the computation of the flux from astrophysical sources and environments, we also provide results for the $\nu$ emissivity, which is the fundamental brick in the flux determination in astrophysical contexts. The flux observed at Earth, $\phi_{\nu}$, is given by the integral along the line-of-sight (l.o.s.) of the $\nu$ emissivity, $\epsilon_{i,j}$, summed over all combinations of incoming CR species $i$ interacting on the ISM components $j$:
$\frac{d^{2}\phi_{\nu}}{d\Omega dE_{\nu}}(E_{\nu},l,b)=\sum\limits_{ij}\int\limits_{\text{l.o.s.}}\!\!\!\!\!\;d\ell\;\;\epsilon_{ij}\Big(\vec{x}(\ell,l,b),E_{\nu}\Big)\,.$ (1)
Here, $\vec{x}$ is the source position, $\ell$ the distance along the l.o.s., while $l$ and $b$ denote the Galactic longitude and latitude, respectively. The flux is expressed as a differential quantity in $\nu$ energy, $E_{\nu}$, and solid angle, $\Omega$. The total flux is computed by integrating over the full solid angle.
The emissivity at a given Galactic location $\vec{x}$ is obtained by convolving the local CR flux at a given position and kinetic energy $T_{i}$, $\phi_{i}(\vec{x},T_{i})$, and the ISM density $n_{\mathrm{ISM},j}(\vec{x})$, with the differential cross section for $\nu$ production, $d\sigma_{ij}(T_{i},E_{\nu})/dE_{\nu}$ for the reaction $i+j\rightarrow\nu+X$:
$\epsilon_{ij}(\vec{x},E_{\nu})=n_{\mathrm{ISM},j}(\vec{x})\int dT_{i}\,\phi_{i}(\vec{x},T_{i})\frac{d\sigma_{ij}}{dE_{\nu}}(T_{i},E_{\nu}).$ (2)
In general, the emissivity varies across the Galaxy, reflecting spatial variations in both the CR flux and the ISM density, and the properties of neutral particles, which are not deflected during propagation.
We deal here with neutrinos produced in hadronic scatterings, mainly $p+p$, therefore arising from the decays of mesons and baryons. In Fig. 1, we illustrate the relevant production channels for $\nu_{\mu}$ and $\nu_{e}$ that have been included in the present analysis. The corresponding antineutrino channels can be obtained by replacing each particle with its antiparticle. The dominant contribution comes from charged pions, which subsequently decay into neutrinos. This channel is examined in detail in Sec. III. The additional channels shown in Fig. 1 are discussed in Sec. IV. Channels contributing less than $0.5\%$ to the total $\nu$ yield are omitted in the figure and neglected in our analysis. Although some of these minor channels are poorly constrained and difficult to quantify precisely, we estimate their combined contribution to be at the level of $\sim 1\%$, which remains below the overall uncertainty budget.
In the following of this section we specify formulae for the pion production case, but they are formally identical for any other meson or baryon directly produced in the $p+p$ collision, as listed in Fig. 1. The $\nu$ production cross section is obtained from the charged-pion one via the convolution:
$\frac{d\sigma_{ij}}{dE_{\nu}}(T_{i},E_{\nu})=\int dT_{\pi^{\pm}}\,\frac{d\sigma_{ij}}{dT_{\pi^{\pm}}}(T_{i},T_{\pi^{\pm}})\;P(T_{\pi^{\pm}},E_{\nu}),$ (3)
where $T_{\pi^{\pm}}$ denotes the kinetic energy of the pion that decays into a $\nu$ with energy $E_{\nu}$. The probability distribution function $P(T_{\pi^{\pm}},E_{\nu})$ for this decay can be computed analytically, and is adopted from Ref. Kelner et al. (2006).
The fully differential production cross section can be expressed in the Lorentz-invariant form:
$\sigma^{(ij)}_{\rm inv}=E_{\pi^{\pm}}\frac{d^{3}\sigma_{ij}}{dp_{\pi^{\pm}}^{3}},$ (4)
Here, $E_{\pi^{\pm}}$ denotes the total energy and $p_{\pi^{\pm}}$ the momentum of the charged pion. This cross section depends on three kinematic variables: the center-of-mass (CM) energy $\sqrt{s}$, the transverse momentum $p_{T}$, and the radial scaling variable $x_{R}$, defined as the ratio of the pion energy to its maximum possible value in the center-of-mass frame, $x_{R}=E_{\pi^{\pm}}^{\ast}/E_{\pi^{\pm}}^{\max\ast}\,.$ The energy-differential cross section in Eq. (3) is obtained by first transforming the kinematic variables from the CM frame to the fixed-target (LAB) frame, and then integrating over the solid angle $\Omega$:
$\frac{d\sigma_{ij}}{dT_{\pi^{\pm}}}(T_{i},T_{\pi^{\pm}})$ $=$ $p_{\pi^{\pm}}\int d\Omega\;\sigma_{{\rm inv}}^{(ij)}(T_{i},T_{\pi^{\pm}},\theta)$
$=$ $2\pi p_{\pi^{\pm}}\int^{+1}_{-1}d(\cos{\theta})\;\sigma_{{\rm inv}}^{(ij)}(T_{i},T_{\pi^{\pm}},\theta)\,$,
where $\theta$ is the angle between the incoming projectile and the outgoing $\pi^{\pm}$ in the LAB frame. In the following, we detail the calculation of the $\nu$ production cross sections from $\pi^{\pm}$ decays, making use of the $\pi^{\pm}$ production cross sections derived in ODDK22.
III $\nu$ from $p+p\rightarrow\pi^{\pm}+X$ collisions
The $\pi^{\pm}$ channel for $\nu$ production corresponds to the first two rows in Fig. 1(left) for $\nu_{\mu}$ and to the first row in Fig. 1(right) for $\nu_{e}$. Given its relevance, it is crucial to obtain precise data covering a wide kinematic phase space for the reaction $p+p\rightarrow\pi^{\pm}+X$. As discussed in ODDK22, measurements of $\sigma_{\rm inv}(p+p\rightarrow\pi^{\pm}+X)$ have been collected by various accelerator and collider experiments, spanning large portions of the kinematic phase space. In particular, NA49 Alt et al. (2005) and NA61 Aduszkiewicz et al. (2017) provide data at low $\sqrt{s}$, while ALICE Aamodt (2011) and CMS Chatrchyan et al. (2012); Sirunyan et al. (2017) cover the high-$\sqrt{s}$ regime. In this work we do not perform a new fit to the particle production data (since no new data have appeared); instead, we use the results for $\sigma_{\rm inv}$ obtained in ODDK22 to start the calculation of $\nu$ yields. The expressions for $\sigma_{\rm inv}$ are given in Eqs. (7)–(10) and the parameters for $\pi^{\pm}$ are listed in Table II of ODDK22. We remark that the fit performed in ODDK22 is based on an analytical formula for $\sigma_{\rm inv}$, directly calibrated on the data and their uncertainties. The highest projectile energy on which we tested our model is $T_{p}=9\times 10^{7}$ GeV, corresponding to $\sqrt{s}=13$ TeV pion data from CMS Sirunyan et al. (2017). Beyond this limit, our model can only be extrapolated. However, care must be taken when treating purely statistical uncertainties at high $\sqrt{s}$. As explained in ODDK22, our model of $\sigma_{\rm inv}$ relies on a fit on NA49 and NA61 $\pi^{\pm}$ data at $\sqrt{s}=17.3$ GeV, which provide broad coverage in both $p_{T}$ and $x_{R}$, thereby constraining in a solid way the dependence of the model on these variables. The scaling of the cross section at higher energies was instead determined by fits to CMS and ALICE data at $\sqrt{s}>900$ GeV (corresponding to $T_{p}=4.3\times 10^{5}$ GeV in the LAB frame). These measurements, however, were performed at mid-rapidity and cover only a limited range in $p_{T}$, with no data available at forward rapidities. Consequently, our model implies an extrapolation of the cross section in the forward-rapidity region at high $\sqrt{s}$. To quantify the impact of this incomplete phase-space coverage, we will compare our results with predictions from other state-of-the-art Monte Carlo models, thereby estimating the overall uncertainties across different cross-section parametrizations.
In Fig. 2, we show the differential cross section for the production of $\nu_{\mu}$, $\bar{\nu}_{\mu}$, $\nu_{e}$, and $\bar{\nu}_{e}$ from $\pi^{\pm}$ decays in $p+p$ collisions. Results are provided for different incident kinetic proton energies $T_{p}$ as a function of $E_{\nu}$, for all $\nu$ flavors. The uncertainties range from $5\%$ to $20\%$ over most of the energy range, except for $E_{\nu}$ values close to $T_{p}$, where statistical errors increase. The most accurate predictions are obtained around $T_{p}\sim 100~\mathrm{GeV}$, corresponding to the energy range covered by NA49 and NA61 data for $\pi^{\pm}$ production. Both $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ receive contributions from $\pi^{+}$ and $\pi^{-}$ decays, as both pion charge can produce a $\nu_{\mu}$. In contrast, $\nu_{e}$ and $\bar{\nu}_{e}$ are primarily produced by $\pi^{+}$ and $\pi^{-}$ decays, respectively. As previously noted in ODDK22, the modeling of $\pi^{-}$ production carries slightly larger uncertainties due to less precise experimental data, leading to $\bar{\nu}_{e}$ having the largest model uncertainty among all $\nu$ flavors.
Figure 2: Differential cross section for the production of $\nu_{\mu}$, $\nu_{e}$ and relative antiparticles from $\pi^{\pm}$ in $p+p$ collisions, computed for different incident kinetic energies of the proton $T_{p}$.
IV Contribution from other production channels and from nuclei
In this section, we present our model for $\nu$ production from all the production channels other than the $\pi^{\pm}$ ones discussed in Sect. III. We also remind the tools for treating scatterings involving nuclei heavier than hydrogen.
The relevant decay channels contributing to $\nu_{\mu}$ production, for which we have an analytical model calibrated on the available data, along with their corresponding branching ratios (in parentheses), are:
  • $K^{+}\rightarrow\mu^{+}\nu_{\mu}$ (63.6%)
  • $K^{+}\rightarrow\pi^{+}\pi^{0}$ (20.7%)
  • $K^{+}\rightarrow\pi^{+}\pi^{+}\pi^{-}$ (5.58%)
  • $K^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}$ (63.6%)
  • $K^{-}\rightarrow\pi^{-}\pi^{0}$ (20.7%)
  • $K^{-}\rightarrow\pi^{-}\pi^{-}\pi^{+}$ (5.58%)
  • $K^{0}_{S}\rightarrow\pi^{+}\pi^{-}$ (69.2%)
  • $\Lambda\rightarrow p\pi^{-}$ (63.9%)
The channels contributing to $\nu_{e}$ production include:
  • $K^{+}\rightarrow\mu^{+}\nu_{\mu}$ (63.6%)
  • $K^{+}\rightarrow\pi^{+}\pi^{0}$ (20.7%)
  • $K^{+}\rightarrow\pi^{+}\pi^{+}\pi^{-}$ (5.58%)
  • $K^{+}\rightarrow\pi^{0}e^{+}\nu_{e}$ (5.1%)
  • $K^{-}\rightarrow\pi^{-}\pi^{-}\pi^{+}$ (5.58%)
  • $K^{0}_{S}\rightarrow\pi^{+}\pi^{-}$ (69.2%)
The corresponding decay channels of the antiparticles contribute to the production of $\bar{\nu}_{\mu}$ and $\bar{\nu}_{e}$.
We include the production cross sections derived in ODDK22. The $\nu$ spectra are calculated assuming that $\pi^{\pm}$ are produced via either a two-body or a three-body decay. In particular, for three-body decays we follow the approach in ODDK22, assuming that each of the three particles carries one-third of the parent’s energy.
The decay time of the $K^{0}_{L}$ meson is $5.1\times 10^{-8}$ s, about 600 times longer than that of $K^{0}_{S}$, which makes detecting $K^{0}_{L}$ particles in accelerator experiments particularly challenging. Moreover, the $K^{0}_{L}$ has different decay channels and branching ratios compared to $K^{0}_{S}$, and contributes to $\nu_{\mu}$ production through:
  • $K^{0}_{L}\rightarrow\pi^{\pm}e^{\mp}\nu_{e}$ ($B_{r}=40.6\%$),
  • $K^{0}_{L}\rightarrow\pi^{\pm}\mu^{\mp}\nu_{\mu}$ ($B_{r}=27.0\%$),
  • $K^{0}_{L}\rightarrow\pi^{+}\pi^{-}\pi^{0}$ ($B_{r}=12.5\%$).
The lack of experimental data prevents an independent parametrization of the production cross section. We therefore use the Pythia event generator to compare the $p_{T}$ and $x_{F}$ dependence of the final $\nu$ spectra from $K^{0}_{S}$ and $K^{0}_{L}$ decays, where $x_{F}=2p_{L}/\sqrt{s}$ and $p_{L}$ is the particle longitudinal momentum. We find that the $p_{T}$ and $x_{F}$ distributions for $\nu_{\mu}$ production are very similar for $K^{0}_{L}$ and $K^{0}_{S}$, differing only in normalization. Considering the different branching ratios for $K^{0}_{S}$ and $K^{0}_{L}$ and their decay chains, one finds that the $\nu_{\mu}$ yield from $K^{0}_{L}$ is about $94\%$ of the $K^{0}_{S}$ one. Similarly, for $\nu_{e}$ the yield from $K^{0}_{L}$ is 1.16 times than from $K^{0}_{S}$. In the following, we therefore assume that the $\nu_{\mu}$ production cross section from $K^{0}_{L}$ can be obtained by rescaling that from $K^{0}_{S}$ by a factor $0.94$, and the $\nu_{e}$ one by a factor $1.16$. By charge symmetry, the same scaling factors are applied to $\bar{\nu}_{\mu}$ and $\bar{\nu}_{e}$. No additional uncertainty is assigned to these numbers, as they follow directly from the well-measured branching ratios of $K^{0}_{L}$ and $K^{0}_{S}$ decays into pions. We then assign to the $K^{0}_{L}$ channel the same uncertainty of the $K^{0}_{S}$ one.
IV.1 Subdominant channels
[…]
Other channels contribute only subdominantly to the yields of $\nu_{\mu}$, $\nu_{e}$, and their respective antiparticles. The $\bar{\Lambda}$, charged $\Sigma$, and $\Xi$ hyperons have typical lifetimes of order $10^{-10}$ s. Their contributions to pion production are usually not included since they decay outside the detector. We therefore include them explicitly in our calculations, while neglecting other minor contributions. Since no experimental data are available, in ODDK22 we estimated the contributions of the $\bar{\Lambda}$, $\Sigma$, and $\Xi$ baryons using the Pythia code Sjöstrand et al. (2015). Specifically, we assume that the contributions of these minor channels to $\nu_{\mu}$ and $\nu_{e}$ are equal to that of $\Lambda$ to $\nu_{\mu}$ and $\bar{\nu}_{e}$, rescaled by a normalization factor $\mathcal{F}$ determined with Pythia. This factor accounts for the production multiplicities of the different hyperons and their branching ratios into channels that can decay into $\pi^{\pm}$. This approach is justified because we have an explicit model for the invariant cross section of $\Lambda$ production, and, given that the masses of $\bar{\Lambda}$, $\Sigma$, and $\Xi$ are similar or identical, we expect their cross sections to share comparable kinematic dependencies.
Figure 4: Comparison among our differential cross sections and the one reported in Kamae et al. (2006) (Kamae), Bhatt et al. (2020) (Bhatt) and Koldobskiy et al. (2021) (AAfrag) as a function of $E_{\nu}$, for incident proton energy $T_{p}=10^{8},10^{6},10^{4}$ and $10^{2}$ GeV.
In Fig. 9 of ODDK22, the $\mathcal{F}$ functions for $e^{+}$ and $e^{-}$ are shown. For $\nu_{\mu}$, the corresponding correction factor is given by the sum of the two functions calculated for $e^{\pm}$, since both $\pi^{+}$ and $\pi^{-}$ contribute to $\nu_{\mu}$ production. In contrast, $\nu_{e}$ production corresponds solely to the $\mathcal{F}_{e^{+}}$ function.
At low energies, $\mathcal{F}$ ranges from 20% to 100% for $\nu_{\mu}$ and from 10% to 50% for $\nu_{e}$, while at high energies it reaches up to 3 for $\nu_{\mu}$ and 2 for $\nu_{e}$. The details of the calculation are provided in ODDK22. The correction factor can vary by as much as 40% depending on the Monte Carlo setup; therefore, we associate a systematic uncertainty of 40% with these channels across all energies.
IV.2 Production from heavier nuclei
For the inclusion of scatterings involving nuclei heavier than hydrogen, either in the CRs or in the ISM, we closely follow the prescriptions derived in ODDK22 for $\pi^{\pm}$. Specifically, when a $\pi^{\pm}$ is produced in collisions between a projectile and a target nucleus with mass numbers $A_{1}$ and $A_{2}$, the $\sigma_{\rm inv}$ formulas reported in Eqs.(7)–(10) of ODDK22 are modified according to the prescriptions in Eqs.(25)–(27) of ODDK22. The parameters in Eq.(26) are taken from the columns 1 and 2 Table V, where the columns for $\pi^{+}$ ($\pi^{-}$) are used to correct $\sigma_{\rm inv}$. The $K^{\pm}$ channel is modified analogously using columns 3 and 4 of Tab. V in ODDK22. For all other channels, we apply a correction function given by the average of the $K^{+}$ and $K^{-}$ cases.
V Results on the $\nu$ production cross section and emissivity
Figure 5: Relative difference between our differential cross section and the one reported in Kamae et al. (2006) (Kamae), Bhatt et al. (2020) (Bhatt) and Koldobskiy et al. (2021) (AAfrag), for incident proton energy $T_{p}=10^{2},10^{4},10^{6}$ and $10^{8}$ GeV for $\nu_{\mu}$. We also show the statistical uncertainty of our model as a gray band.
We have now all the tools to compute the total differential cross section $d\sigma/dE_{\nu}$ for the inclusive production of neutrinos in inelastic $p+p$ collisions. The result is obtained by summing all contributions from $\pi^{\pm}$ and the subdominant channels, as discussed in Sections III and IV. These cross sections constitute the main result of our paper and are shown in Fig. 3 for a representative proton energy of $T_{p}=10^{6}$ GeV, for all flavors of neutrinos and antineutrinos. The $\pi^{\pm}$ channels dominate at all $\nu$ energies and for all flavors. All kaons contribute about 20% of the pions and 15% of the total, while all other subdominant channels scrape the 1-2% of the total. We have verified that these patterns are maintained across different proton energies. The gray curve and shaded band show the total $d\sigma/dE_{\nu}$ and its $1\sigma$ uncertainty, respectively. The resulting uncertainty spans from $6\%$ to $20\%$ depending on $E_{\nu}$, and is primarily driven by the modeling of the $\pi^{\pm}$ cross section, with larger uncertainties for $\bar{\nu}_{e}$ due to the dominant production of $\pi^{-}$, as explained in Section III. We remind that our cross-section model is validated at high $\sqrt{s}$ by a fit to ALICE and CMS data probing the midrapidity region only. This ensures that we capture the correct dependence on $\sqrt{s}$ of the pion multiplicity, which is predominantly produced at midrapidity. However, the spectral shape of our cross section—depending on $p_{T}$ and $x_{R}$—relies on extrapolation at high $\sqrt{s}$.
We compare our results, obtained from data-driven analytical approach, with state-of-the-art Monte Carlo codes. In Fig. 4 we show $d\sigma/dE_{\nu}$ for different neutrino flavors and results from Monte Carlo codes as a function of neutrino energy and for various incident proton energies. Fig. 5 (shown for $\nu_{\mu}$ only) zooms on the relative difference between our model and the Monte Carlo predictions, with the statistical uncertainty of our model indicated by a gray band. As a general remark, our cross sections are in good agreement with the predictions by Bhatt, with discrepancies of the order of $10\%$, except for $E_{\nu}\sim T_{p}$, where the cross sections are nevertheless close to zero. The AAfrag model is also generally consistent with our results, particularly at low energies where deviations remain below $10\%$, while at higher energies differences increase to approximately $20\%$.
Figure 6: The $\nu$ emissivity is computed for $p+p$, He+p+p, p+p+He, He+He and CNO+p+p scatterings for the different neutrino and antineutrino flavors. The gray line is the sum of all contributions (see text for details). Each prediction is plotted with the relevant uncertainty due to the production cross sections derived in this paper. In each bottom panel, the relative uncertainty to the total $\epsilon(E_{\nu})$ is displayed. For comparison, we show the results by Kamae et al. (2006) (Kamae), Bhatt et al. (2020) (Bhatt) and Koldobskiy et al. (2021) (AAfrag) in the $p+p$ channel.
In contrast, the Kamae et al. model shows deviations up to 50%, particularly at low energies, and displays significantly different behaviours across the flavors, with the exception of $\nu_{e}$, which remains close to the other models. We do not report Kamae model for $T_{p}=10^{6},10^{8}$ GeV, since they do not provide the tables for these energies. This discrepancy could be particularly relevant for the observation of a significant Galactic diffuse neutrino emission made by the IceCube Neutrino Observatory in Abbasi et al. (2023), where their template model is based on Kamae’s cross sections.
In order to assess the impact of our new cross sections, we compute the emissivity for the different $\nu$ flavors and their corresponding antiparticles as in Eq. 2, assuming a constant $n_{\rm ISM}$ ($n_{\rm H}=0.9~{\rm cm}^{-3}$ and $n_{\rm He}=0.1~{\rm cm}^{-3}$) and incident CR spectra that are independent of Galactic position. The CR spectra are obtained from a fit to the AMS-02 Aguilar et al. (2021) and CALET data Adriani et al. (2022, 2023). In Fig. 6, we show $\epsilon(E_{\nu})$ as a function of $E_{\nu}$ for $p+p$, He+p+p, p+p+He, He+He, and CNO+p+p interactions, along with their sum. We plot two curves for the total emissivity, both shown in gray (solid and dot-dashed): one is obtained by extrapolating the high-energy behavior of the CR spectrum from fits to AMS-02 and CALET data (solid line), and the other assuming an exponential cutoff (dot-dashed) in the Galactic CR spectrum at 1 PeV for protons (scaling with the charge Z of each element, as the maximum energy is expected to be rigidity dependent). The presence or absence of such a cutoff at PeV energies leads to noticeable differences in the predicted emissivity already at $10^{4}$ GeV, consistent with the fact that predictions at a given $E_{\nu}$ depend in a non-trivial way on the integration over incident CR energies at least two orders of magnitude higher. Therefore, the accuracy of any neutrino production cross sections depends on the proton energy or, equivalently, the beam energy at accelerators, well beyond the desired $E_{\nu}$.
Each prediction provides the corresponding uncertainty from the production cross section. The relative uncertainty on the total $\epsilon(E_{\nu})$ is shown in the bottom panels. As expected, the dominant contribution comes from $p+p$ interactions. However, scatterings involving He collectively yield a similar emissivity, with the He+p+p channel becoming comparable to $p+p$ at very high energies for all $\nu$ flavors, due to the different high-energy behaviour hinted at by CALET measurements Adriani et al. (2023). The decrease of the pp/He CR ratio with energy explains why the predicted emissivities from $p+p$ and He+p+p interactions converge at high energies, given that the cross section for the latter is roughly 3.15 times larger than that of the $p+p$ channel. This plot is intended as an illustrative example to highlight the impact of our cross-section model once convolved with typical astrophysical source spectra.
The statistical uncertainty on $\epsilon(E_{\nu})$ due to hadronic production cross sections is approximately $5\%$ for $E_{\nu}\lesssim 10^{3}$ GeV, and increases to $10\%$ for $E_{\nu}\gtrsim 10^{6}$ GeV. For comparison, we show the results in dashed lines from Kamae et al. (2006) (Kamae), Koldobskiy et al. (2021) (AAfrag), and Bhatt et al. (2020) (Bhatt) for the $p+p$ channel. The models by Kamae and Bhatt do not provide full coverage across the proton and neutrino energy range required for this plot, and are therefore cut accordingly. There is excellent agreement at low energies ($E_{\nu}\lesssim 100$ GeV) among our model, AAfrag, and Bhatt. Overall, our results and Bhatt’s (orange line) agree within $\sim 10\%$, while AAfrag shows deviations up to $\sim 20\%$. Conversely, Kamae exhibits significant discrepancies below $10^{3}$ GeV (except for $\nu_{e}$), while remaining broadly consistent with the other models between $10^{3}$ GeV and $10^{4}$ GeV. As a general trend, our new cross sections predict an emissivity that, at least for the dominant $p+p$ channel, is a bit higher that the Monte Carlo state of the art.
VI Discussion and conclusions
Hadronic collisions represent a primary mechanism for generating high-energy neutrinos in astrophysical environments. The neutrino production is dominated by the decay of charged pions, in turn produced by the inelastic scattering of CR nuclei with the target material. A precise modeling of the neutrino production cross section from hadronic interactions is crucial for interpreting current neutrino data.
In this paper, we present a new evaluation of the neutrino production cross section from $p+p$ collisions, building on our previous analysis of the production cross section for $\pi^{\pm}$, $K^{\pm}$, and minor baryonic and mesonic channels that contribute at least at the 0.5% level. Cross sections for scatterings involving nuclei heavier than protons are also derived. The novelty of our approach, compared to the state of the art that relies on Monte Carlo predictions, is that we shape the $\sigma_{\rm inv}$ analytically, and directly fit the model to the available data. We work with neutrino energies from $10$ GeV to $10^{7}$ GeV, and correspondingly to incident proton (nuclei) energies from $10$ GeV to $10^{9}$ GeV (GeV/n). Our results include a realistic and conservative estimate of the uncertainties affecting the differential cross section $d\sigma/dE_{\nu}$, understood as the sum over all production channels. The uncertainty is estimated to be $5\%$ for $E_{\nu}\leq 1000$ GeV, increasing to $10\%$ at higher energies in the TeV range and reaching 20% at $E_{\nu}\simeq 10^{7}$ GeV.
We compare our results to cross sections from state-of-the-art Monte Carlo codes, finding good agreement with Bhatt et al. (2020) and Koldobskiy et al. (2021), but not with Kamae et al. (2006), which exhibits significant discrepancies with our predictions.
To improve the accuracy of current results, new collider data are required. In particular, measurements of the Lorentz-invariant cross section for $\pi^{\pm}$ production at forward rapidities are essential. Recent efforts at the LHC to explore previously unmeasured regions of phase space – such as those by LHCf Adriani et al. (2016), FASER Abraham et al. (2024), and SND@LHC Acampora et al. (2024) – could also be leveraged to provide measurements relevant to neutrino production, for reactions of interest in astrophysical studies Maurin et al. (2025). It would also be important to perform similar measurements using a He target. Our results could also be relevant for proton beam-dump and fixed-target searches for light dark matter, where neutrino-induced neutral-current interactions constitute leading backgrounds Abreu et al. (2023); Albanese et al. (2023); Batell et al. (2022); Ahdida et al. (2022).
We provide numerical tables for the energy-differential cross sections $d\sigma/dE_{\nu}$ as a function of $E_{\nu}$ from $10$ GeV to $10^{7}$ GeV, and for incident proton (nuclei) energies from $10$ GeV to $10^{9}$ GeV (GeV/n). A script to read the tables is also included. The material is available at https://github.com/lucaorusa/neutrino_cross_section.
Acknowledgments
LO thanks Julien Dörner for inspiring discussions. LO acknowledges the support of the Multimessenger Plasma Physics Center (MPPC), NSF grant PHY2206607. FD thanks the Department of Theoretical Physics of CERN, where part of this work for carried on. MDM and FD acknowledge support from the research grant TAsP (Theoretical Astroparticle Physics) funded by Istituto Nazionale di Fisica Nucleare (INFN), MDM from the Italian Ministry of University and Research (MUR), PRIN 2022 “EXSKALIBUR – Euclid-Cross-SKA: Likelihood Inference Building for Universe’s Research”, Grant No. 20222BBYB9, CUP I53D23000610 0006, and from the European Union – Next Generation EU, and FD from the Research grants The Dark Universe: A Synergic Multimessenger Approach, No. 2017X7X85K, funded by the Miur.